OpenStudy (anonymous):

Let T be in S(sub7). Suppose T^3= (1,2,3,4,5,6,7), a 7-cycle. Determine what permutation T represents and justify your answer.

OpenStudy (kinggeorge):

Well let's think about it. We know that T^3 has order 7, so T itself must have order dividing 3*7=21. Since it obviously can't have order 1 or 3, T must have order 7. So let's write the permutation as follows.\[1\mapsto?\mapsto?\mapsto2\]\[2\mapsto?\mapsto?\mapsto3\]\[3\mapsto?\mapsto?\mapsto4\]\[4\mapsto?\mapsto?\mapsto5\]\[5\mapsto?\mapsto?\mapsto6\]\[6\mapsto?\mapsto?\mapsto7\]\[7\mapsto?\mapsto?\mapsto1\]Each time we travel across the arrow, we're applying T. Since we've determined that T has order 7, let's follow the arrow 7 times for each number, and wee what we get.\[1\mapsto6\mapsto?\mapsto2\]\[2\mapsto7\mapsto?\mapsto3\]\[3\mapsto1\mapsto?\mapsto4\]\[4\mapsto2\mapsto?\mapsto5\]\[5\mapsto3\mapsto?\mapsto6\]\[6\mapsto4\mapsto?\mapsto7\]\[7\mapsto5\mapsto?\mapsto1\]But this all we need. We've exactly described T as \[T=(1642753)\]

OpenStudy (kinggeorge):

Did the process make sense?

OpenStudy (anonymous):

I'm not sure I follow what the areas are doing. We just went over this in class today and it made NO sense at all

OpenStudy (kinggeorge):

The arrows? They just represent applying the permutation \(T\). So permutation the (1234567) could be written as \[1\mapsto2\]\[2\mapsto3\]\[3\mapsto4\]\[4\mapsto5\]\[5\mapsto6\]\[6\mapsto7\]\[7\mapsto1\]In the work I showed above, I was using the arrows as an alternate notation for \(T\). So by moving across the arrows 3 times, we apply T three times, which is the same as applying \(T^3\).

OpenStudy (anonymous):

ok thank you! but how did you get (1642753)? I get that we are applying T 3 times, but how did you get that answer. I'm sorry I just really don't get this kind of thing

OpenStudy (kinggeorge):

You understand why T^7 is the identity correct?

OpenStudy (anonymous):

....not really. I know that the order of S7 is 7! and that the order of T is 7!/2 which is 504.....I think that's right

OpenStudy (kinggeorge):

Alright, so the order of \(S_7\) is 7!. The order of \(T\) must therefore divide \(7!\). But you know that the order of \(T^3\) is 7 (since it's a 7-cycle). Thus, \((T^3)^7=T^{21}=e\). Still following me?

OpenStudy (anonymous):

yes! this makes much more sense than class

OpenStudy (kinggeorge):

Excellent. Then since \(T^{21}=e\), you know that if \(T^r=e\), and \(r\) is minimal, then \(r\) must divide 21. The only divisors of 21 are 1,3,7,21, so we only need to see if \(T\) can have order 1,3,7 or 21. You still understand this? Oh, and btw, I just thought of a simpler way to get find \(T\), but we still need this information.

OpenStudy (anonymous):

ok go on. I'm still with you

OpenStudy (kinggeorge):

Great. If \(T\) has order 1, then \(T=e\), and \(T^3=e\neq(1234567)\). So the order of \(T\) is not 1. If \(T\) has order 3, then \(T^3=e\), so it's the same thing if it had order 1. Now, this statement is the only part where I'll do a little hand-waving. There are no elements in \(S_7\) of order 21. You can check this yourself if you know how to compute the order of a general element in \(S_n\). Otherwise, just take my word for it. Thus, \(T\) must have order 7. Still understanding what I'm doing?

OpenStudy (anonymous):

yes

OpenStudy (kinggeorge):

Now is when I'll start to do things a bit differently. Since the order of \(T\) is 7, we know that \(T^7=e\), so \(T^6=T^{-1}\). But we're given \(T^3\). So we can easily find \(T^6=(T^3)^2\). Can you tell me what \(T^6\) is?

OpenStudy (anonymous):

....the inverse? I don't know

OpenStudy (kinggeorge):

Think of \(T^6\) as \((T^3)^2\) and not the inverse.

OpenStudy (anonymous):

so it's what we had originally, we just permutate....twice?

OpenStudy (kinggeorge):

Precisely.

OpenStudy (anonymous):

....so....I don't follow what you are trying to tell me from that.

OpenStudy (kinggeorge):

If \(T^3=(1234567)\), what is \((T^3)^2=(1234567)(1234567)\)?

OpenStudy (anonymous):

is this mod 7?

OpenStudy (kinggeorge):

I guess you could think of it like that. It shouldn't matter though.

OpenStudy (anonymous):

well am i multiplying them or not? This is the part that he didn't cover in class. would everything just shift one? so then its (7123456)?

OpenStudy (kinggeorge):

Alright. Let's review how to multiply two permutations with a small example. Suppose the permutation we have is \((13425)\), and we want to find \((13425)(13425)\). Let's look at where 1 is sent to. We apply (13425) to 1, and we get 3. Now if we apply (13425) again, 3 gets sent to 4. We ended with 4, so let's see what happens when we apply (13425) to 4. Well, it gets sent to 2. Applying (13425) again we see that 2 is sent to 5. We do it again with 5. Using the same process, 5 gets sent to 1, which is then sent to 3. With three, it gets sent to 4, and then 2. Finally, 2 gets sent first to 5, which then is sent 1. Thus, we can conclude that (13425)(13425)=(14532).

OpenStudy (kinggeorge):

Let me know if you have questions about that.

OpenStudy (anonymous):

so far I have (1723456)......

OpenStudy (kinggeorge):

Take a step back. Do you understand how multiplication works with permutations?

OpenStudy (kinggeorge):

For example, can you find (123)(234)?

OpenStudy (anonymous):

no we've never done it before.... I'm assuming that 1 goes to 2 3 goes to 3 and 3 goes to 4

OpenStudy (anonymous):

so maybe (143)......I really don't know.

OpenStudy (kinggeorge):

No. First look at 1. If we apply (234) to 1, nothing happens. But then we apply (123) which sends 1 to 2. Now we look at 2. (234) sends 2 to 3. But then we apply (123) to 3, to get 1. Since we just found where 1 was sent, we know that (123)(234) has the 2-cycle (12) Next, look at 3. This immediately gets sent to 4, but (123) doesn't change 4. Finally, look at 4. (234) sends 4 to 2, and then (123) sends 2 to 3. So we also have the 2-cycle (34). Therefore, (123)(234)=(12)(34).

OpenStudy (anonymous):

ohhhhhhhhhhhh ok. but we only want one 7-cycle a

OpenStudy (anonymous):

as our answer right?

OpenStudy (kinggeorge):

In the example we're working with, we do want to end up with a 7-cycle. So let's try one more time. What's (1234567)(1234567)?

OpenStudy (anonymous):

ok.... 1 goes to 2, 2 goes to 3, 3 goes to 4, 4 goes to 5, 5 goes to 6, 6 goes to 7.....7 then goes to 1 and then 2.....from there I'm not sure

OpenStudy (kinggeorge):

I'm not doing a great job at how to multiply permutations. You might want to check up with your teacher about how to do it, or read through these links. They give a pretty decent explanation. http://math.stackexchange.com/questions/31763/multiplication-on-permutation-group-written-in-cyclic-notation http://www.mymathforum.com/viewtopic.php?f=13&t=10019 But for now, lets keep working with (1234567) step by step.

OpenStudy (kinggeorge):

First. Let's look at 1. After applying (1234567) twice, what does 1 get sent to?

OpenStudy (anonymous):

...3?

OpenStudy (anonymous):

because 1 goes to 2 which goes to 3?

OpenStudy (kinggeorge):

Perfect. What happens to 3 if you apply (1234567) twice?

OpenStudy (anonymous):

it goes to 5?

OpenStudy (kinggeorge):

Right again! Where does 5 end up? What about 7?

OpenStudy (anonymous):

5 goes to 7.....and 7 goes to 2

OpenStudy (kinggeorge):

Perfect again. Continuing this, 2 goes to 4, 4 goes to 6, and 6 goes to 1. Putting this all together, we get that (1234567)(1234567)=(1357246).

OpenStudy (anonymous):

HOW!? That's where I'm confused. we just said (6712345) so how are you getting (1357246)

OpenStudy (anonymous):

oh wait......wait i get it. holy cow i see it now cause 1 goes to 3 and 3 goes to 5....etc???????

OpenStudy (kinggeorge):

That's precisely it. Gotta love those A-HA moments :)

OpenStudy (anonymous):

haha so now what? we have T^6..... I need T

OpenStudy (kinggeorge):

Well we know from our discussion way above that \(T^6=T^{-1}\). I have three ways to find \(T\). A difficult way, a slightly confusing way, and a magical way :P Which way would you prefer?

OpenStudy (anonymous):

um....magic lol. Unicorns and rainbows and everything.

OpenStudy (kinggeorge):

Magical way here we go. Suppose you have a \(k\)-cycle \((i_1\,i_2\,...\,i_k)\). Then its inverse is \((i_1\,i_k\,i_{k-1}\,...\,i_3\,i_2)\).

OpenStudy (anonymous):

so (1357246) goes to (1642753)?

OpenStudy (kinggeorge):

Yup. So that's \(T\).

OpenStudy (anonymous):

HOLY CRAP!

OpenStudy (anonymous):

You are amazing! thank you thank you THANK YOU!

OpenStudy (anonymous):

That is the most amazing thing....ever. I love math when I finally understand it. Thanks KingGeorge!!!!

OpenStudy (kinggeorge):

You're welcome. And btw, the slightly confusing would have basically proved the magical way. I'll let you work out the details for yourself if you want to. The difficult way would have been raising \(T^{-1}\) to the 6th power. Which still isn't that bad, but it's not great.